How do you solve these problems by using the
quadratic formula x² = 6x - 4 and
x² + 16x + 60 = 5

Answer by Edwin McCravy(20055)

How do you solve these problems by using the 
quadratic formula x² = 6x - 4 and 
x² + 16x + 60 = 5

                    x² = 6x - 4

Get 0 on the right by adding -6x + 4 to both sides


           x² - 6x + 4 = 0

Use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

where a = 1; b = -6; c = 4

                      _____________ 
             -(-6) ± Ö(-6)²-4(1)(4)
        x = ————————————————————————
                     2(1) 
                  _____ 
             6 ± Ö36-16
        x = ————————————
                 2

                  __ 
             6 ± Ö20
        x = —————————
                2

                  ___ 
             6 ± Ö4·5
        x = ——————————
                2 

                   _
             6 ± 2Ö5
        x = —————————
                2

                     _
             6     2Ö5
        x = ——— ± —————
             2      2
                 _
        x = 3 ± Ö5 
                      _
Using the +, x = 3 + Ö5, which
is one solution and equals about 5.236067977
                      _ 
Using the -, x = 3 - Ö5, which
is the other solution and equals about
.583592135

==============================

                   x² + 16x + 60 = 5
                    
Get 0 on the right by adding -5 to both sides


                  x² + 16x + 55 = 0

Use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

where a = 1; b = 16; c = 55

                      ______________ 
             -(16) ± Ö(16)²-4(1)(55)
        x = ——————————————————————————
                     2(1) 

                    _______ 
             -16 ± Ö256-220
        x = ————————————————
                    2

                    __ 
             -16 ± Ö36
        x = —————————————
                  2

                   
             -16 ± 6
        x = —————————
                2 

            -16     6
        x = ———— ± ———
             2      2 

        x = -8 ± 3

Using the +, x = -8 + 3, which
is one solution and equals -5

Using the -, x = -8 - 3, which
is the other solution and equals -11

Edwin